# ChuLo: Chunk-Level Key Information Representation for Long Document Processing

### v3

@@ 3.4 Chunk Representation Training
In this final step, we train a Transformer-based model using our keyphrase-enhanced chunk representations to effectively incorporate the core semantic content of the document. We selected BERT-base according to Table 8. By emphasizing key information in the chunk embeddings, we ensure that the model can focus on the most relevant aspects of the text, thereby improving its ability to handle long document inputs without losing critical context. To achieve this, we leverage a Transformer-based backbone model, which is used to initialize the weights of the chunk attention module, as illustrated in Figure 1. This chunk attention module is designed to capture the intricate contextual relationships among chunks while retaining the influence of keyphrases. By doing so, the module can better understand local and global semantic patterns, enabling the model to perform robustly across various long document tasks. The chunk embeddings are processed through the chunk attention module to produce refined contextual representations, which are then fed into a classification head to generate the final predictions. Our framework, ChuLo, is adaptable to any transformer-based architecture, from pretrained to large language models, making it versatile for tasks involving long document understanding. Through integrating keyphrase-enhanced chunk representations, the model achieves superior performance in both document classification and token-level tasks, highlighting the effectiveness of our approach in leveraging semantic information to tackle the challenges associated with long document processing.
@@ 4 Experiments set-up
We evaluate ChuLo on document and token classification tasks, highlighting our motivation for including both types. While document classification primarily requires global contextual understanding, token classification tasks test the model’s ability to retain and utilize detailed token-level information within long documents. We compare it with BERT (Kenton & Toutanova, 2019) and BERT variants (Park et al., 2022), Longformer (Beltagy et al., 2020), ToBERT (Pappagari et al., 2019), CogLTX (Ding et al., 2020), ChunkBERT (Jaiswal & Milios, 2023), and instructions with LLMs, GPT4o [^1] https://openai.com/index/hello-gpt-4o/ and Gemini1.5pro [^2] https://deepmind.google/technologies/gemini/pro/. Baselines Details are listed in Appendix A.2.
@@
Datasets: We conduct experiments using three(3) datasets for long document classification and two(2) for long document token classification. For document classification, we use HP(Kiesel et al., 2019), LUN (Rashkin et al., 2017), and Eurlex57k (Chalkidis et al., 2019). These datasets vary in average document length from 707 to 1147 tokens, enabling us to assess performance on documents of different lengths and complexities. Further details on dataset statistics and splits are available in Appendix A.4. 1) HP (Hyperpartisan News Detection): evaluates the classification of news articles based on hyperpartisan argumentation (Kiesel et al., 2019). We use the ‘byarticle’ version, which contains 238 hyperpartisan and 407 non-hyperpartisan articles. The same train-test split as in (Beltagy et al., 2020) is adopted. 2) LUN Used for unreliable news source classification, this dataset includes 17,250 articles from satire, propaganda, and hoaxes (Rashkin et al., 2017). Our goal is to predict the source type for each article. 3) Eurlex57k A multi-label classification dataset consisting of 47,000 English EU legislative documents with 4,271 EUROVOC concepts. We also include a simulated Inverted-Eurlex57k version, where the header and recitals are moved to the end, compelling the model to read the entire document for key information. For token classification, we use GUM and CoNLL-2012 for Named Entity Recognition (NER) tasks: 1) GUM (Georgetown University Multilayer) is a richly annotated collection of 235 documents across various genres such as academic texts, news, fiction, and interviews (Lin & Zeldes, 2021). GUM’s various linguistic styles and structures make it an excellent benchmark for assessing token-level understanding in lengthy documents, ensuring that the model captures complex entity relationships over extended contexts. 2) CoNLL-2012 dataset Originally designed for coreference resolution, this dataset is adapted for NER in our work (Pradhan et al., 2013). We convert the data to a document-level format and select the top-k longest documents in each split, emphasizing the model’s ability to understand and process extended text sequences for token classification tasks.
@@
Metrics and Implementation: For the HP and LUN datasets, we use accuracy as the evaluation metric, while for Eurlex57k, Inverted Eurlex57k, GUM, and CoNLL-2012, we adopt micro F1. These metrics are chosen to maintain consistency with prior work and facilitate direct comparison. Regarding implementation, we provide key details here, with the complete setup in Appendix A.5. We use CrossEntropy loss for training on the Hyperpartisan, LUN, CoNLL and GUM datasets, and Binary CrossEntropy loss for the Eurlex57k and Inverted Eurlex57k datasets. All models are optimized using the AdamW optimizer, and training employs early stopping based on the respective validation metric, with a patience threshold set to 10 epochs. A learning rate search is conducted for each experiment to ensure optimal model performance for comparison. Top-n value is set to 15333We tested with different n values, but 15 was generally better in most datasets.
@@ 5 Results
5.1 Document classification performance
We evaluate the effectiveness of our ChuLo by comparing it with fine-tuned PLMs and previously published baselines (Park et al., 2022; Jaiswal & Milios, 2023) on several benchmark datasets: HP, LUN, EURLEX57K, and Inverted EURLEX57K. The comparative results are summarized in Table 1, with input configurations provided in Table 2 and detailed descriptions available in Appendix A.3. Our method demonstrates clear superiority on three of the four datasets, achieving a significant improvement of 6.43% accuracy on the LUN dataset compared to the second-best model, BERT. This marked improvement presents ChuLo’s ability to capture comprehensive document context through its keyphrase-based chunk representation, despite using only 512 input embeddings. The results suggest that our method effectively mitigates the limitations of traditional truncation and chunking strategies by preserving critical semantic information, which contributes to higher classification accuracy. On the EURLEX57K and Inverted EURLEX57K datasets, ChuLo achieves consistent performance gains over baselines, further validating its capability to handle long documents efficiently. In these datasets, which have hierarchical labels and require understanding complex semantic structures, our model benefits from enhanced chunk representations that emphasize key content. This allows ChuLo to capture document semantics better, even when compared to models that can process larger input lengths. While our model delivers competitive results on the HP dataset, it trails behind Longformer by a slight margin of 0.0031 in accuracy. This marginal difference corresponds to only one additional correctly classified instance out of a total of 65 test samples.
Model HP LUN EURLEX57K I-EURLEX57K
BERT (Kenton & Toutanova, 2019) 0.9200 0.5797 0.7309 0.7053
ToBERT (Pappagari et al., 2019) 0.8954 0.3697 0.6757 0.6731
CogLTX (Ding et al., 2020) 0.9477 - 0.7013 0.7080
Longformer (Beltagy et al., 2020) 0.9569 0.5552 0.5453 0.5647
BERT+TextRank (Park et al., 2022) 0.9115 0.4880 0.7287 0.7130
BERT+Random (Park et al., 2022) 0.8923 0.3015 0.7322 0.7147
ChunkBERT (Jaiswal & Milios, 2023) 0.9300 - 0.6494 0.6294
@@ 5.1 Document classification performance
Table 1: Document classification Result. Following previous work, we use accuracy for HP and LUN, and micro F1 for other datasets. Results for LUN are obtained by our own experiment based on provided baseline codes and methods, while baseline results for the other datasets are from previous work(Park et al., 2022; Jaiswal & Milios, 2023). The best performance for each dataset is bolded while the second best is underscored, and we can see that our final model, a BERT-based backbone, generally outperforms other baselines across all datasets by achieving the best or second-best.
@@
BERT (Kenton & Toutanova, 2019) F-512 tokens
ToBERT (Pappagari et al., 2019) All
CogLTX (Ding et al., 2020) S-512 tokens
Longformer (Beltagy et al., 2020) F-4096 tokens
BERT+TextRank (Park et al., 2022) F-512 + S-512 tokens
BERT+Random (Park et al., 2022) F-512 + S-512 tokens
ChunkBERT (Jaiswal & Milios, 2023) All
@@
Interestingly, for the other datasets, Longformer underperforms compared to models like BERT variants or CogLTX, which use the first 512 tokens and focus on selecting key sentences. This observation indicates that unfiltered additional content can introduce noise, negatively impacting classification accuracy. In contrast, ChuLo expands the input content and strategically emphasizes key semantic elements during chunk representation. This approach mitigates noise interference, ensuring that only the most relevant information is retained and highlighted. ChuLo achieves superior performance by balancing content and semantic emphasis, outperforming other models. Overall, the results confirm that ChuLo consistently outperforms standard PLM baselines and existing chunking methods in long document classification tasks. Its ability to retain and emphasize key semantic content, while efficiently handling long inputs, makes it a robust solution for various document classification challenges. The subsequent sections delve deeper into the impact of our chunk representation strategy and discuss its contributions to improving document classification performance.
5.2 Document Classification Performance in longer documents
To further validate the robustness of our model, we evaluate its classification performance across various document length ranges, with a particular focus on longer documents. For this analysis, we consider the documents with more than 1024 tokens and more than 2048 tokens in the test set. To provide a fair comparison, we use Longformer based on its original code and hyperparameters as described in Table 1 and off-the-shelf LLMs, GPT4o and Gemini1.5 pro. As shown in Table 3, our model consistently outperforms others on longer documents in the LUN dataset. Specifically, for documents exceeding 2,048 tokens, ChuLo maintains a higher accuracy compared to all baselines, demonstrating its capacity to handle lengthy inputs effectively. This performance gain can be attributed to our chunk representation’s emphasis on keyphrases, which preserves crucial semantic content even when document length increases. On the HP dataset, ChuLo and Longformer achieve perfect accuracy (1.0) for documents longer than 2,048 tokens. However, for shorter documents (more than 1,024 tokens), ChuLo surpasses Longformer. This improvement is likely due to our chunk representation strategy, which selectively highlights key content rather than averaging information across the entire document. As a result, ChuLo maintains high semantic fidelity, leading to better overall performance even with condensed text inputs.
@@ 5.2 Document Classification Performance in longer documents
We also benchmarked against newly released LLMs, GPT-4o and Gemini 1.5 Pro, using longer document inputs for both the LUN and HP datasets. On LUN, GPT-4o achieved an accuracy of 0.7143 and Gemini 1.5 Pro scored 0.6531, both surpassing Longformer. However, ChuLo achieved the highest accuracy of 0.7959, showcasing its superiority in handling long documents with diverse content. On the HP dataset, GPT-4o (0.8889) and Gemini 1.5 Pro (0.7778) performed worse than Longformer and ChuLo, both of which achieved a perfect accuracy of 1.0 on the longer documents. This highlights ChuLo’s robustness and consistency in classifying documents with varying length, even compared to advanced language models. The prompt and response samples are in Appendix A.6 and A.7. Overall, these results demonstrate that ChuLo not only outperforms standard PLM baselines and chunking methods on long documents but also remains competitive against the latest large language models. By prioritizing key semantic elements and effectively managing document length, ChuLo maintains stable performance across varying input lengths.
@@
5.3 Token Classification Performance
To further demonstrate the effectiveness of our chunk representation method, we evaluated it on a token-level classification task—specifically, Named Entity Recognition (NER) using long documents. We compared our model against two state-of-the-art long-document pre-trained models, Longformer (Beltagy et al., 2020) and BigBird (Zaheer et al., 2020), as well as newly released large language models, GPT-4o and Gemini 1.5 Pro. As shown in Table 4, our model consistently outperforms both Longformer and BigBird and LLM models on the NER tasks, particularly on the CoNLL, where document lengths often exceed the input limitations of these baseline models. To leverage the broader context captured by our chunk representation, we integrate a BERT-decoder module that utilizes the enhanced chunk embeddings to predict token labels more accurately. This configuration allows the model to maintain a global understanding of the document while focusing on the local dependencies necessary for precise token classification. The superior performance of our ChuLo on the CoNLL demonstrates the significance of retaining global contextual information when dealing with long documents. All baselines struggle with these longer inputs due to their limited capacity for handling extensive sequences. In contrast, our method’s ability to encode the entire document’s context through keyphrase-based chunk representations enables it to achieve higher accuracy in recognizing and classifying named entities. This is particularly evident in cases where long-distance dependencies and contextual nuances play a critical role in determining the correct labels. Overall, the results indicate that our model’s chunk representation not only enhances performance on document-level classification tasks but also proves highly effective for token-level tasks such as NER. By retaining global context while emphasizing key semantic content, our method enables more accurate token classification, validating its application in downstream tasks that require detailed and comprehensive understanding of long document tokens.
5.4 Token Classification Performance in longer documents
We further analyze the NER performance across different document length ranges. As presented in Table 5(a) and Table 5(b), we report the number of documents exceeding specific length thresholds and their corresponding performance metrics. On the CoNLL, as document lengths exceed the maximum input capacities of Longformer and BigBird, both models exhibit significant performance drops to 31.56% and 31.45%, respectively. In contrast, our model experiences a minimal decrease of 1.28%, showcasing its resilience and effectiveness in handling long sequences. For the GUM, where all document lengths are within the acceptable range for these models, performance remains stable across all models, with our approach consistently achieving the best results.
CoNLL Entire dataset (20) Longer than 2048 (17) Longer than 4096(6) Longer than 8192 (2)
@@ 5.4 Token Classification Performance in longer documents
GUM Entire dataset (26) - Longer than 512 Longer than 1000(8) Longer than 1042 (6)
@@
Table 5: NER results for comparison on documents of different lengths. Values of brackets indicate the # of documents for each document. The best performance (Micro F1) is bolded and the second best is underscored, and our model consistently outperforms all the baselines for each document set.
@@
Figure 2: Comparison of performance in different length ranges for CoNLL and GUM datasets. Values of brackets includes the min and max length of each dataset

### v4

@@ 3.4 Chunk Representation Training
In this final step, we train a Transformer-based model using our keyphrase-enhanced chunk representations to effectively incorporate the core semantic content of the document. We selected BERT-base according to Table 14. By emphasizing key information in the chunk embeddings, we ensure that the model can focus on the most relevant aspects of the text, thereby improving its ability to handle long document inputs without losing critical context. We leverage a Transformer-based backbone model, which is used to initialize the weights of the chunk attention module, as illustrated in Figure 1. This chunk attention module is designed to capture the intricate contextual relationships among chunks while retaining the influence of keyphrases. By doing so, the module can better understand local and global semantic patterns, enabling the model to perform robustly across various long document tasks. The chunk embeddings are processed through the chunk attention module to produce refined contextual representations, which are then fed into a classification head to generate the final predictions. Our framework, ChuLo, is adaptable to any transformer-based architecture, from pretrained to large language models, making it versatile for tasks involving long document understanding. Through integrating keyphrase-enhanced chunk representations, the model achieves superior performance in both document classification and token-level tasks, highlighting the effectiveness of our approach in leveraging semantic information to tackle the challenges associated with long document processing.
@@ 4 Experiments Set-up
We evaluate ChuLo on document and token classification tasks to highlight our motivation. While document classification primarily requires global contextual understanding, token classification tasks test the model’s ability to retain and utilize detailed token-level information within long documents. We compare it with BERT (Kenton and Toutanova, 2019) and BERT variants (Park et al., 2022), Longformer (Beltagy et al., 2020), ToBERT (Pappagari et al., 2019), CogLTX (Ding et al., 2020), ChunkBERT (Jaiswal and Milios, 2023), and instructions with LLMs, GPT4o [^3] https://openai.com/index/hello-gpt-4o/ and Gemini1.5pro [^4] https://deepmind.google/technologies/gemini/pro/. Baselines Details are listed in Appendix A.3.
@@
Datasets: We conduct experiments using three(3) datasets for long document classification and two(2) for long document token classification. For document classification, we use HP(Kiesel et al., 2019), LUN (Rashkin et al., 2017), and Eurlex57k (Chalkidis et al., 2019). These datasets vary in average document length from 707 to 1147 tokens, enabling us to assess performance on documents of different lengths and complexities. Further details on dataset statistics and splits are available in Appendix A.5. 1) HP (Hyperpartisan News Detection) evaluates the classification of news articles based on hyperpartisan argumentation (Kiesel et al., 2019). We use the ‘byarticle’ version, which contains 238 hyperpartisan and 407 non-hyperpartisan articles. The same train-test split as in (Beltagy et al., 2020) is adopted. 2) LUN uses for unreliable news source classification, this dataset includes 17,250 articles from satire, propaganda, and hoaxes (Rashkin et al., 2017). Our goal is to predict the source type for each article. 3) Eurlex57k consists of 47,000 English EU legislative documents with 4,271 EUROVOC concepts. We also include a simulated Inverted-Eurlex57k version, where the header and recitals are moved to the end, compelling the model to read the entire document for key information. For token classification, we use GUM and CoNLL-2012 for Named Entity Recognition (NER) tasks: 1) GUM (Georgetown University Multilayer) is a richly annotated collection of 235 documents across various genres such as academic texts, news, fiction, and interviews (Lin and Zeldes, 2021). GUM’s various linguistic styles and structures make it an excellent benchmark for assessing token-level understanding in lengthy documents, ensuring that the model captures complex entity relationships over extended contexts. 2) CoNLL-2012 originally designed for coreference resolution, and is adapted for NER in our work (Pradhan et al., 2013). We convert the data to a document-level format and select the top-k longest documents in each split, emphasizing the model’s ability to process extended text sequences for token classification tasks.
@@
Metrics and Implementation: For the HP and LUN datasets, we use accuracy as the evaluation metric, while for Eurlex57k, Inverted Eurlex57k, GUM, and CoNLL-2012, we adopt micro F1. These metrics are chosen to maintain consistency with prior work and facilitate direct comparison. Regarding implementation, we provide key details here, with the complete setup in Appendix A.6. We use CrossEntropy loss for training on the Hyperpartisan, LUN, CoNLL and GUM datasets, and Binary CrossEntropy loss for the Eurlex57k and Inverted Eurlex57k datasets. All models are optimized using the AdamW optimizer, and training employs early stopping based on the respective validation metric, with a patience threshold set to 10 epochs. A learning rate search is conducted for each experiment to ensure optimal model performance for comparison. Top-n value is set to 15555We tested with different n values, but 15 was generally better in most datasets.
@@ 5 Results
5.1 Document Classification
We evaluate the effectiveness of our ChuLo by comparing it with fine-tuned PLMs and previously published baselines (Park et al., 2022; Jaiswal and Milios, 2023) on several benchmark datasets: HP, LUN, EURLEX57K, and Inverted EURLEX57K. The comparative results are summarized in Table 1, with input configurations provided in Table 2 and detailed descriptions available in Appendix A.4. Our method demonstrates clear superiority on three of the four datasets, achieving a significant improvement of 6.43% accuracy on the LUN dataset compared to the second-best model, BERT. This marked improvement presents ChuLo’s ability to capture comprehensive document context through its keyphrase-based chunk representation, despite using only 512 input embeddings. The results suggest that our method effectively mitigates the limitations of traditional truncation and chunking strategies by preserving critical semantic information, which contributes to higher classification accuracy. On the EURLEX57K and Inverted EURLEX57K datasets, ChuLo achieves consistent performance gains over baselines, further validating its capability to handle long documents efficiently. In these datasets, which have hierarchical labels and require understanding complex semantic structures, our model benefits from enhanced chunk representations that emphasize key content. This allows ChuLo to capture document semantics better, even when compared to models that can process larger input lengths. While our model delivers competitive results on the HP dataset, it trails behind Longformer by a slight margin of 0.0031 in accuracy. This marginal difference corresponds to only one additional correctly classified instance out of a total of 65 test samples.
Model HP LUN EUR I-EUR
BERT 0.9200 0.5797 0.7309 0.7053
ToBERT 0.8954 0.3697 0.6757 0.6731
CogLTX 0.9477 - 0.7013 0.7080
Longformer 0.9569 0.5552 0.5453 0.5647
BERT+TextRank 0.9115 0.4880 0.7287 0.7130
BERT+Random 0.8923 0.3015 0.7322 0.7147
ChunkBERT 0.9300 - 0.6494 0.6294
@@ 5.1 Document Classification
Table 1: Document classification Result. Following previous work, we use accuracy for HP and LUN, and micro F1 for other datasets. Results for LUN are obtained by our own experiment based on provided baseline codes and methods, while baseline results for the other datasets are from previous work(Park et al., 2022; Jaiswal and Milios, 2023). are the BERT variants proposed by Park et al. (2022). The best performance for each dataset is bolded while the second best is underscored, and we can see that our final model, a BERT-based backbone, generally outperforms other baselines across all datasets by achieving the best or second-best.
@@
BERT F-512 tokens
ToBERT All
CogLTX S-512 tokens
Longformer F-4096 tokens
BERT+TextRank F-512 + S-512 tokens
BERT+Random F-512 + S-512 tokens
ChunkBERT All
@@
Interestingly, for the other datasets, Longformer underperforms compared to models like BERT variants or CogLTX, which use the first 512 tokens and focus on selecting key sentences. This observation indicates that unfiltered additional content can introduce noise, negatively impacting classification accuracy. In contrast, ChuLo expands the input content and strategically emphasizes key semantic elements during chunk representation. This approach mitigates noise interference, ensuring that only the most relevant information is retained and highlighted. Overall, the results confirm that ChuLo consistently outperforms standard PLM baselines and existing chunking methods in long document classification tasks. Its ability to retain and emphasize key semantic content, while efficiently handling long inputs, makes it a robust solution for various document classification challenges.
5.2 Longer Document Classification
To further validate the robustness of our model, we evaluate its classification performance across various document length ranges, with a particular focus on longer documents. For this analysis, we consider the documents with more than 1024 tokens and more than 2048 tokens in the test set. We use Longformer and off-the-shelf LLMs, GPT4o and Gemini1.5 pro for comparison. As shown in Table 3, our model consistently outperforms others on longer documents in the LUN dataset. Specifically, for documents exceeding 2,048 tokens, ChuLo maintains a higher accuracy compared to all baselines, demonstrating its capacity to handle lengthy inputs effectively. This performance gain can be attributed to our chunk representation’s emphasis on keyphrases, which preserves crucial semantic content even when document length increases. On the HP dataset, ChuLo and Longformer achieve perfect accuracy (1.0) for documents longer than 2,048 tokens. However, for shorter documents (more than 1,024 tokens), ChuLo surpasses Longformer. This improvement is likely due to our chunk representation strategy, which selectively highlights key content rather than averaging information across the entire document. As a result, ChuLo maintains high semantic fidelity, leading to better overall performance even with condensed text inputs.
@@ 5.2 Longer Document Classification
We also benchmarked against newly released LLMs, GPT-4o and Gemini 1.5 Pro, using longer document inputs for both the LUN and HP datasets. On LUN, GPT-4o achieved an accuracy of 0.7143 and Gemini 1.5 Pro scored 0.6531, both surpassing Longformer. However, ChuLo achieved the highest accuracy of 0.7959, showcasing its superiority in handling long documents with diverse content. On the HP dataset, GPT-4o (0.8889) and Gemini 1.5 Pro (0.7778) performed worse than Longformer and ChuLo, both of which achieved a perfect accuracy of 1.0 on the longer documents. This highlights ChuLo’s robustness and consistency in classifying documents with varying length, even compared to advanced language models. The prompt and response samples are in Section 5.5 and A.8. Overall, these results demonstrate that ChuLo not only outperforms standard PLM baselines and chunking methods on long documents but also remains competitive against the latest large language models. By prioritizing key semantic elements and managing document length, ChuLo maintains stable performance across varying input lengths.
@@
5.3 Token Classification
To further demonstrate the effectiveness of our chunk representation method, we evaluated it on a token-level classification task—specifically, Named Entity Recognition (NER) using long documents. We compared our model against two state-of-the-art long-document pre-trained models, Longformer (Beltagy et al., 2020) and BigBird (Zaheer et al., 2020), as well as newly released large language models, GPT-4o and Gemini 1.5 Pro. As shown in Table 4, our model consistently outperforms Longformer, BigBird and LLM models on the NER tasks, particularly on the CoNLL, where document lengths often exceed the input limitations of these baseline models. To leverage the broader context captured by our chunk representation, we integrate a BERT-decoder module that utilizes the enhanced chunk embeddings to predict token labels more accurately. This configuration allows the model to maintain a global understanding of the document while focusing on the local dependencies necessary for precise token classification. All baselines struggle with these longer inputs due to their limited capacity for handling extensive sequences. In contrast, our method’s ability to encode the entire document’s context through keyphrase-based chunk representations enables it to achieve higher accuracy in recognizing and classifying named entities. This is particularly evident in cases where long-distance dependencies and contextual nuances play a critical role in determining the correct labels. Overall, the results indicate that our model’s chunk representation not only enhances performance on document-level classification tasks but also proves highly effective for token-level tasks such as NER, validating its application in downstream tasks that require a detailed and comprehensive understanding of long document tokens.
CoNLL ALL (20) > 2048 (17) > 4096(6) > 8192 (2)
@@ 5.3 Token Classification
GUM ALL (26) - > 512 > 1000(8) > 1042(6)
@@
Table 5: NER results for comparison on documents of different lengths. >number represents the documents longer than the number, with the values in brackets indicating the corresponding counts for the documents. The best performance (Micro F1) is bolded and the second best is underscored, and our model consistently outperforms all the baselines for each document set.
@@
Figure 2: Comparison of performance in different length ranges for CoNLL and GUM datasets. Values of brackets includes the min and max length of each dataset.
5.4 Token Classification in Longer Documents